Status of calibrations for 2003 data

From: Bjorn H Samset (bjornhs@rcf2.rhic.bnl.gov)
Date: Tue Sep 09 2003 - 08:10:29 EDT

  • Next message: Kris Hagel: "brat update on afs"
    Hi dev'ils.
    
    Last week in Copenhagen, we had a mini-workshop between NBI, Oslo and
    Bergen to kickstart the process of making the rest of the calibrations
    needed for the 2003 data. Here is a short update on what we've been doing
    so far and what we're planning.
    
    TPC calibrations
    ----------------
    TPC calibrations have been made and commited to the db for the following
    periods:
    dAu: 8307-8632 (later part of the run, where inel disc. was OK)
    pp: 8675-9030 (entire 2003 run)
    So far we've done one calib. per setting (i.e. angle and mag. field) but
    since we calibrate both spectrometers every time most long settings have
    had several calibrations checked in. (E.g. if the FS setting was changed
    several times while the MRS remained constant)
    
    Local tracking files
    --------------------
    ...have been made for all the above mentioned runs, and can be found here:
    /brahms/data07/data/run03/[dau/pp]/200/rXXXX/ltr/ltr*.root
    
    Naming scheme
    -------------
    You'll note that the above naming scheme is a bit different from what
    we've been used to. This is the result of a discussion that has been going
    for some time on how to name the official files - the main points are:
    * Official files will be kept under this kind of structure:
    /brahms/dataXY/data/run/system/energy/r1234/kind/kind123456seq123v12.root
    where
    'kind' is either ltr, gtr or dst (and some more...)
    'v12' is a version number
    * The filenames are all lowercase
    * There is a script (run2ltr.py) that moves the output from local tracking
    to the above structure
    This is of course not nesecearily the final word, but we're testing it now
    to see if it works the way we intended...
    
    Matching offsets
    ----------------
    ...have been made for the dAu runs and are currently being made for pp.
    They will be put in an official location once they've been checked a bit.
    
    Global tracking
    ---------------
    ...is done for dAu, not yet for pp. This also needs to be checked a bit,
    and will then be put in the structure showed above.
    
    Pedestal calibrations ++ for globals and hodoscopes
    ------------------------------------------------
    * Hiro is currently doing what remains of pedestal and slewing
    calibrations for TD1, MrsT0, BB and INEL.
    * Claus has updated the RICH calib software so that it can (hopefully)
    handle the double-peaked pedestals we saw for part of the run. He's doing
    ped. cal's now.
    * Claus has also, as you've seen, started doing ZCD ped. calibrations
    where possible.
    * Once all this is in, we'll make a set of TOF calibrations (H1, H2, TOFW)
    - this is a job for the Danish-Norwegian CalibSquad, and for anyone else
    who wants to contribute. Hands?
    
    DSTs
    ----
    ...will then be made and put in the structure shown above.
    
    Reco- and calibscripts
    ----------------------
    ...will be placed in a similar structure to the data to keep track of
    what was used. Currenly, a beginning of such a structure is maintained
    under
    ~bramreco/bhs/reco/run03/etc...
    and
    ~bramreco/bhs/calib/run03/etc...
    More will be added here as we go, and the scripts we finally use will be
    put in a more official and permanent location.
    
    Other stuff
    -----------
    In addition to this, there is of course
    * the DCs, which are being calibrated in Krakow
    * the Mult, which I guess is more or less working (from Steves analysis
    note)
    * the fieldmaps, where I'd like a status report from someone in the know
    * C4, which is being worked on in Kansas
    * TOFW2, which I know naught about...
    * ...and probably more that I've currently forgotten.
    
    I'm sure many steps will need redoing a number of times before we have
    something we trust, but we have to start somewhere and this is hopefully
    it. We'll keep you posted as to what's happening on our side, and please
    contribute your knowledge, wisdom and calibrations as applicable :-)
    
    P.S: As we move towards Quark Matter and more analysis, we really need to
    clean up our data disks. We've started a discussion with Flemming on this
    - a suggestion for a Disk Use Policy and requests for data moving/removal
    will be sent out in not too long a time...
    
    Ping :-)
    
    --
    Bjorn H. Samset                           Phone: 22856465/92051998
    PhD student, heavy ion physics            Adr:   Schouterrassen 6
    Inst. of Physics, University of Oslo             0573 Oslo
                                  \|/
    ----------------------------> -*- <-----------------------------
                                  /|\
    


    This archive was generated by hypermail 2.1.5 : Tue Sep 09 2003 - 08:11:39 EDT